# EF-Calib: Spatiotemporal Calibration of Event- and Frame-Based Cameras Using Continuous-Time Trajectories

### v1

@@ V-B3 Time Offset Calibration Test
8.1633 2.8632 4.6642 2.6372 1.2161 1.4511
8.0019 2.8187 3.8794 2.7859 0.4974 1.4322
0.9141 5.6910 0.6322 0.3239 0.3707 0.3065
0.5336 0.1984 0.6572 0.3127 0.3638 0.2911
V-C Ablation Study
To thoroughly analyze and validate the performance and functionality of each module within EF-Calib, we conducted an ablation study. Specifically, we scrutinized three modules: piece-wise trajectories, temporal calibration, and feature refinement. Table III demonstrates the impact of the introduction of these three modules on the calibration error of the EF-Calib extrinsic parameters. As can be seen from Table III, the introduction of all three modules can greatly improve the extrinsic parameter calibration accuracy.
@@ VI Conclusion
In this letter, we propose a spatiotemporal calibration framework called EF-Calib, aiming to achieve joint calibration of intrinsic parameters, extrinsic parameters, and time offset for event and frame-based cameras. First, we design a novel calibration pattern that accommodates the heterogeneous nature of event and frame-based representations, incorporating both circles and crosspoints to facilitate simultaneous recognition by both camera types. A corresponding event-based recognition algorithm is developed to ensure robust and accurate feature recovery using this pattern. Additionally, to manage the asynchronous characteristics of the event stream, we introduce a piece-wise B-spline to continuously represent the pose trajectory of the event camera. Finally, we provide the analytic Jacobian of the error term and implement the joint calibration of intrinsic, extrinsic, and time offset for both camera types. Experimental results demonstrate that EF-Calib outperforms current state-of-the-art methods in intrinsic parameter estimation while also achieving high accuracy in extrinsic parameter and time offset estimation. These results demonstrate the spatiotemporal calibration capabilities of EF-Calib and lay a robust foundation for the fusion of event and frame.

### v2

@@ V-C Ablation Studies
8.163 2.863 4.664 2.637 1.216 1.451
8.002 2.819 3.879 2.786 0.497 1.432
0.914 5.691 0.632 0.324 0.371 0.307
0.534 0.198 0.657 0.313 0.364 0.291
TABLE IV: Ablation Study of Optimization Strategies
Two-stage Separation Trial 1 Trial 2 Trial 3
$e_{t}$ $e_{r}$ $e_{t}$ $e_{r}$ $e_{t}$ $e_{r}$
11.026 0.628 1.429 0.351 1.600 0.390
0.552 0.211 0.697 0.320 0.383 0.289
0.548 0.213 0.696 0.320 0.382 0.289
To thoroughly analyze and validate the performance of each module within EF-Calib, as well as the optimization strategy employed, we conducted two ablation studies. The first study focused on three key modules: piece-wise trajectories, temporal calibration, and feature refinement. Table III shows the impact of these modules on the calibration error of EF-Calib’s extrinsic parameters. The Baseline algorithm, as listed in the top row, represents the entire event camera motion trajectory with a single continuous B-spline curve, assumes no time offset between cameras, and does not apply feature refinement. As Table III illustrates, incorporating these three modules significantly enhances calibration accuracy and ensures convergence to the correct solution.
The second ablation study compared different optimization strategies. Table IV demonstrates that our strategy not only improves calibration accuracy over single-stage joint optimization but also significantly reduces computational load while maintaining accuracy comparable to the incremental two-stage optimization. This efficiency is attributed to the first stage’s effective refinement of parameters to near-optimal values, making the EF-Calib optimization strategy both highly accurate and efficient.
@@ VI Conclusion
In this letter, we propose a novel calibration framework called EF-Calib, aiming to achieve spatiotemporal calibration of intrinsic parameters, extrinsic parameters, and time offset for event and frame-based cameras. Experimental results demonstrate that EF-Calib outperforms current state-of-the-art methods in intrinsic parameter estimation while also achieving high accuracy in extrinsic parameter and time offset estimation. These results demonstrate the spatiotemporal calibration capabilities of EF-Calib and lay a robust foundation for the fusion of event and frame.